OpenStudy (mayankdevnani):

The time period of a satellite is 5 hours. If the separation between the earth and the satellite is increased to 4 times what will be the new time period? a) 20 hrs b) 10 hrs c) 80 hrs d) 40 hrs

OpenStudy (anonymous):

20h i guess

OpenStudy (mayankdevnani):

it is absolutely wrong sis? @ShikhaDessai

OpenStudy (mayankdevnani):

@amistre64 @Callisto @pratu043 @inkyvoyd @eyust707 @rajathsbhat @Aperogalics @sirm3d @Shadowys plz help me guys:)

OpenStudy (anonymous):

i'm not sure........never solved this kind of a problem before

OpenStudy (anonymous):

kepler's third law bud. \[\Large \frac{T_1^{2}}{T_2^{2}}=\frac{r_1^{3}}{r_2^{3}}\]

OpenStudy (mayankdevnani):

ok...then

OpenStudy (mayankdevnani):

why should we use kepler's 3 law, why should'nt we use 5*4=20 hrs

OpenStudy (anonymous):

umm because the universe obeys the law of gravity?

OpenStudy (anonymous):

let height be h first and whoose time period is 5hrs h---->4h by keplers third law